# Measuring the Dunkelflaute: How (not) to analyze variable renewable energy shortage

### v1

@@ 2.3 Comparability and the choice of threshold
Comparability issues. Research interest is increasingly moving beyond single technologies at particular regions (e.g., cannon_using_2015 ; potisomporn_spatial_2022 ; potisomporn_evaluating_2023 ; abdelaziz_assessing_2024 ; ohlendorf_frequency_2020 ) towards a characterization of VRE drought patterns of portfolios of different renewable generation technologies for multiple regions (e.g., raynaud_energy_2018 ; rinaldi_wind_2021 ; kies_critical_2021 ; kapica_potential_2024 ). For this, enabling comparability across VRE technologies and regions when analyzing VRE drought events is imperative.
@@
Differences in generation potentials: single technologies. Previous research draws on absolute, exogenously set thresholds for availability factors, e.g., 0.01, 0.02, 0.05, or 0.10 leahy_persistence_2013 ; patlakas_low_2017 ; kaspar_climatological_2019 ; ohlendorf_frequency_2020 ; kies_critical_2021 ; potisomporn_extreme_2024 . Yet, such analyses face compatibility issues when analyzing systems with different technologies and/or across several regions. This is because applying absolute thresholds on normalized availability factor time series implies assuming uniform generation capacity (e.g., in megawatts) across all compared systems. However, these systems may have very different annual generation potentials (e.g., in megawatt hours) depending on the meteorological conditions of the location and the VRE technology. VRE availability duration curves, sorting availability factor time series in descending order, highlight the discrepancies in generation potentials between different technologies and regions (Figure 4a).
@@
Figure 4: Annual generation potentials of onshore wind power, solar PV, and portfolios with different mixes of these technologies in Germany and Spain in 2010, visualized as the area below the availability duration curves. The data is retrieved from the Pan-European Climate Database de_felice_entso-e_2022 .
@@
Differences in generation potentials: technology portfolios. The same applies to energy systems transitioning towards variable renewables. Here, the optimal capacity mix of VRE technologies depends on their respective generation potentials as well as cost parameters, and may vary substantially across regions. This results in diverse generation potentials for such VRE portfolios based on the location and the capacity mix (Figure 4b).
@@
Challenges with absolute thresholds: FLH bias. Full-load hours (FLH) serve as a commonly used metric for comparing annual average generation (or generation potential before potential curtailment). This metric indicates how long a generator (or technology aggregate) would operate at full capacity to provide its annual energy (potential). An analysis based on a uniform, exogenously set threshold for each compared system does not account for discrepancies in annual generation potentials. Consequently, such analyses (e.g., kaspar_climatological_2019 ; kies_critical_2021 ; ohba_climatology_2022 ; mayer_probabilistic_2023 ) are biased by a FLH effect, which puts their validity into question.
@@
Threshold scaling relative to maximum availability factor. As a remedy, the drought threshold can be scaled to account for FLH differences between VRE technologies and locations. Threshold scaling relative to different time series attributes can be found in the literature. One option is to scale thresholds relative to their system-specific maximum availability factor breyer_reflecting_2022 . [^6] Technically, in this study, thresholds scale with the maximum generation. Discarding potential curtailment, this is equivalent to scaling with the maximum availability factor.. Focusing on the VRE availability of a single point in time (i.e., the peak point of the duration curves in Figure 4), this approach ignores general availability patterns over the course of a year. Consequently, it does not mitigate differences in FLH.
@@
Thresholds scaling relative to the distribution of availability factors. Another option is to use a percentile of the availability time series distribution as threshold fleig_global_2006 ; cannon_using_2015 ; plain_accounting_2019 . Duration curves are a graphical representation of availability factor distributions (Figure 4). A percentile refers to the availability factor below a certain fraction of the population of time steps falls and depends on the shape of the distribution, which may vary across technologies and regions. Unlike scaling relative to the maximum availability factor, this method recognizes all VRE availability states. However, the percentile approach does not fully adjust for FLH differences.
@@
Thresholds scaling relative to mean availability factor. An effective option for this is to align the drought threshold with each system’s mean availability factor $\bar{\phi}^{VRE}$. This method inherently scales with the system’s FLH and, therefore, enables comparability across technologies or technology portfolios, regions, and, in the case of comparing multiple years, time raynaud_energy_2018 ; rinaldi_wind_2021 ; gangopadhyay_role_2022 ; otero_copula-based_2022 ; kapica_potential_2024 .
@@
Effective threshold ranges. A priori, a meaningful parameterization of the fraction $\alpha$ of the mean availability factor $\bar{\phi}^{VRE}$ for drought identification is not straightforward. Previous studies use various fractions of the mean yevyevich_objective_1967 . While lower thresholds, e.g., $\tau=0.1\bar{\phi}^{VRE}$, find very severe but potentially brief and isolated VRE drought events, higher fractions closer to $\bar{\phi}^{VRE}$ may result in identifying extended periods of below-average availability rather than true VRE droughts. Relevant fractions likely lie between these extremes. It therefore appears useful to explore a range of thresholds, $\tau=\alpha*\bar{\phi}^{VRE}$, for example with $\alpha$ ranging between 0.1 and 0.9. Obtained events reflect different degrees of low VRE availability, which may be relevant for different types of flexibility options in renewable energy systems.
@@ 2.4 Technological and spatial scope of analysis
VRE drought analyses scopes. The scope of a VRE drought analysis focusing on VRE supply may differ in terms of technological and spatial scope. Each of these dimensions demands specific assumptions, weighting, as well as comparability requirements.
Portfolio time series in multi-technology analyses. In contrast to single-technology drought analyses, according investigations of a VRE portfolio require a composite time series that combines the availability factors of different VRE technologies. It can be computed as the weighted average of the technology-specific time series. Weighting factors correspond to the technology shares in the assumed overall capacity mix. Considering the temporal and spatial complementarity of wind and solar power weschenfelder_review_2020 ; pedruzzi_review_2023 , such a composite time series allows for within-portfolio balancing, which reflects the generation dynamics of power sectors with high shares of variable renewables. This is in line with several related studies raynaud_energy_2018 ; otero_characterizing_2022 ; ruhnau_storage_2022 ; francois_statistical_2022 . Another approach excludes or limits such balancing by applying a definition of compound VRE drought events as periods with simultaneously arising drought conditions of multiple VRE technologies rinaldi_wind_2021 ; mayer_probabilistic_2023 . This discards VRE complementarity, hence potentially overestimating droughts characteristics of a VRE portfolio.
Interconnection states in multi-regional analysis. For multi-regional settings, drought analyses using VRE availability time series can be based on two extreme scenarios regarding regional interconnection [^7] Here we abstract from within-region grid congestion.: i) perfect interconnection, which abstracts from grid limitations (often referred to as the “copper plate” assumption), and ii) complete isolation, where regions are treated as “island systems” without electricity exchange. When comparing VRE droughts of such island systems, FLH-adjusted threshold scaling is required as discussed above. Realistic patterns of cross-regional electricity exchange fall between these extremes and typically necessitate more sophisticated energy system modeling beyond simple VRE time series analysis.
Cross-regional time series. Similar to the technology portfolio case, multi-regional settings with assumed perfect interconnection require a composite time series that collectively represents all regions. It can be generated as the weighted average of respective regional time series. The weights are assigned based on the relative share of each region’s capacity within the overall system. Unless multiple cross-regional composite time series are to be compared, no FLH-adjustment of the threshold is required. For VRE portfolios across regions, weighting factors need to reflect both within-regional and across-regional shares in the capacity mixes.
@@ 3.1 Positive residual load event definition
Motivation: kalte Dunkelflauten. In addition to focusing exclusively on VRE supply, research interest on the dynamics between VRE droughts and periods of high electricity demand is intensifying. This becomes more relevant as many countries plan to electrify their heat, transport, and industry sectors, and respective additional loads enter the power sector. Notably during the heating season, electrified heating appliances such as heat pumps increase electricity demand with characteristic profiles bloess_power–heat_2018 ; ruhnau_heating_2020 ; roth_flexible_2023 . Compound high-demand and VRE drought events, sometimes also referred to as “cold dark doldrums” or, in German, “kalte Dunkelflauten” are emerging as a challenge for decarbonized energy systems that rely on VRE.
@@
Positive & negative residual load events. PRL events, characterized by high demand and a shortage of VRE supply, indicate the need for firm capacity from thermal generators, or flexibility provided by electricity storage, imports, or demand-side measures schill_electricity_2020 ; zerrahn2018 ; goke_how_2023 . In contrast, negative residual load events represent VRE surplus generation. This surplus can be integrated into the power sector through electricity storage or exports to other areas, or used for electrified loads from coupled sectors, such as heat provision or transport brown_synergies_2018 . Alternatively, it can be curtailed schill_residual_2014 ; zerrahn2018 . Note that the latter is not an option for PRL periods, which makes them more challenging to deal with than negative residual loads. Accurately quantifying both positive and negative residual load events is relevant for capacity planning of weather-resilient energy systems.
@@
PRL characteristics. Key PRL characteristics, e.g., duration and energy deficits, can be analyzed using residual load time series $\phi_{t}^{RL}$ (Figure 5). While this can be done analogously to VRE drought analysis, the energy deficit of PRL events can directly be associated with the need for energy that has to be provided by, for instance, long-duration storage. Importantly, the maximum capacity need can be inferred from peak residual load (compare left-most value of the residual load duration curve in Figure 5) and does not require a dedicated analysis of PRL events.
@@
Implicit threshold parameterization. Unlike VRE drought analysis, there is no need for deliberate threshold parameterization to identify PRL events. Instead, PRL events are simply obtained as periods above zero (gray line in Figure 5). In previous PRL analysis other approaches such as percentile-based thresholds were used otero_copula-based_2022 . However, positive residual load directly relates to the need for energy from firm generators, storage, or imports. Hence, from the perspective of a power sector, diverting from thresholds aligned with the zero line appears counter-intuitive.
[…]
3.2 Identification methods
@@
Identification methods: CAZ & MAZ. We adapt the VRE droughts identification methods discussed in Section 2 to PRL events. Unless stated otherwise, the algorithmic routine is equivalent but recalibrated to identify periods where residual load $\phi_{t}^{RL}$, or its average $\Phi_{t}^{RL}$, is above zero, rather than below a designated threshold. Analogously to CBT and MBT, we denote Constantly-Above-Zero (CAZ) and Mean-Above-Zero (MAZ) as:
$\psi_{t}^{CAZ}=$$\begin{cases}1&\text{if $\phi_{t}^{RL}>0\ \ \forall\ \ t$}\\ 0&\text{if $\phi_{t}^{RL}\leq 0\ \ \forall\ \ t$}\end{cases}$ $\psi_{t}^{FMAZ}=$$\begin{cases}1&\text{if $\Phi_{t}^{RL}(\omega)>0\ \ \forall\ \ t$}\\ 0&\text{if $\Phi_{t}^{RL}(\omega)\leq 0\ \ \forall\ \ t$}\end{cases}$ (5)
@@
Identification methods: FMAZ & VMAZ. Accordingly, the fixed-duration MAZ (FMAZ) and variable-duration MAZ (VMAZ) concepts emulate the FMBT and VMBT approaches.
@@
Duration CAZ, FMAZ, VMAZ.. Analogously to VRE drought analysis, the PRL event duration $\delta^{CAZ}$ and $\delta^{FMAZ}$ corresponds to the number of consecutive time steps with $\psi_{t}^{CAZ}=1$ and $\psi_{t}^{FMAZ}=1$, respectively. The duration of VMAZ event equals the iteration-specific averaging interval $\delta^{VMAZ}=\omega^{VMAZ}$.
@@
Energy deficit CAZ, FMAZ, VMAZ. The VMAZ energy deficit is near zero and hence insignificant. With $t_{k}$ and $t_{l}$ as the start and end points of a PRL event, the CAZ and FMAZ event’s energy deficit (green and orange areas in Figure 5) are defined as:
$\epsilon^{CAZ}=$ $\sum_{t_{k}}^{t_{l}}\phi_{t}^{RL}$ $\epsilon^{FMAZ}=$ $\sum_{t_{k}}^{t_{l}}\Phi_{t}^{RL}$ $\forall\ \ t_{k}\leq t\leq t_{l}$ (6)
@@
Limitations of CAZ, FMAZ, and VMAZ. The limitations associated with CBT, FMBT, and VMBT discussed in Section 2 also extend to their PRL counterparts. The CAZ method, while directly linking event duration, its energy deficit, and its temporal occurrence to for example the required generation and dispatch period of a capacity-constrained firm generator, ignores brief periods of negative residual load adjacent to PRL events (compare the two CAZ events on the left in Figure 5). This can lead to underestimating the need for energy-constrained flexibility options, such as long-duration electricity storage. As for FMAZ, the parameterization of the averaging interval $\omega^{MAZ}$ lacks a directly attributable power sector reference and remains arbitrary. As identified FMAZ events may substantially change with varying $\omega^{MAZ}$, they are ambiguous. Further, their interpretation is not straightforward. In general, longer averaging periods may lead to an underestimation of actual PRL challenges as averaging already assumes temporal balancing, for instance, in real-world power systems carried out by electricity storage or other flexibility options. The VMAZ approach, in contrast, obtains unique events. Yet, its relevance remains limited in PRL analysis as it only allows for evaluating PRL event duration but not energy deficit, which converges to zero by design.
@@
Identification method: SPA. The SPA identifies unique PRL while obtaining a meaningful energy deficit. Unlike for VRE droughts, SPA does not require setting an arbitrary threshold for PRL detection, precluding a respective shortcoming in the case of VRE drought identification. SPA allows for brief periods with negative residual load and detects all PRL events within a single iteration. Consequently, we regard it as particularly well-suited for identifying PRL events. Analogously to VRE droughts, its cumulative energy deficit is denoted as:
$\epsilon_{t}^{SPA}=\begin{cases}\epsilon^{SPA}_{t-1}+\phi_{t}^{RL}&\text{if $\epsilon^{SPA}_{t-1}+\phi_{t}^{RL}>0\ \ \forall\ \ t$}\\ 0&\text{if $\epsilon^{SPA}_{t-1}+\phi_{t}^{RL}\leq 0\ \ \forall\ \ t$}\end{cases}$ (7)
@@
SPA duration and recovery period. The SPA event duration $\delta^{SPA}$ refers to the period beginning when the cumulative energy deficit of a single SPA event becomes positive to its global maximum. The SPA method further allows for quantifying a recovery period $\sigma^{SPC}$. Originating from the hydrology field duckstein_engineering_1987 , this period represents the balancing time required for the surplus VRE energy from intermediate and subsequent negative residual load events to offset the energy deficit of a PRL event. It corresponds to the period following the peak cumulative energy deficit $\epsilon_{t}^{SPA}$ to the point in time where this deficit returns to zero $\epsilon_{t}^{SPC}=0$ (Figure 5).
[…]
$\phi_{t}^{RL,adj}=\frac{\phi_{t}^{RL}}{\eta^{LDS}}\ \ \forall\ \ \phi_{t}^{RL}<0$ (8)
[…]

### v2

@@ 2.3 Comparability requirements and the choice of threshold
Research interest is increasingly moving beyond single technologies at particular regions (e.g., [24, 25, 26, 29, 32]) towards a characterization of VRE drought patterns of portfolios of different renewable generation technologies for multiple regions (e.g., [19, 18, 34, 35]). For this, enabling comparability across VRE technologies and regions when analyzing VRE drought events is imperative.
@@
Previous research draws on absolute, exogenously set thresholds for availability factors, e.g., 0.01, 0.02, 0.05, or 0.10 [30, 31, 38, 32, 34, 17]. Yet, such analyses face compatibility issues when analyzing systems with different technologies and/or across several regions. This is because applying absolute thresholds on normalized availability factor time series implies assuming uniform generation capacity (e.g., in megawatts) across all compared systems. However, these systems may have very different annual generation potentials (e.g., in megawatt hours) depending on the meteorological conditions of the location and the VRE technology. VRE availability duration curves, sorting availability factor time series in descending order, highlight the discrepancies in generation potentials between different technologies and regions (Figure 4a).
@@
Figure 4: Annual generation potentials of onshore wind power, solar PV, and portfolios with different mixes of these technologies in Germany and Spain in 2010, visualized as the area below the availability duration curves. The data is retrieved from the Pan-European Climate Database [56].
@@
The same applies to energy systems transitioning towards variable renewables. Here, the optimal capacity mix of VRE technologies depends on their respective generation potentials as well as cost parameters, and may vary substantially across regions. This results in diverse generation potentials for such VRE portfolios based on the location and the capacity mix (Figure 4b).
@@
Full-load hours (FLH) serve as a commonly used metric for comparing annual average generation (or generation potential before potential curtailment). This metric indicates how long a generator (or technology aggregate) would operate at full capacity to provide its annual energy (potential). An analysis based on a uniform, exogenously set threshold for each compared system does not account for discrepancies in annual generation potentials. Consequently, such analyses (e.g., [38, 34, 41, 40]) are biased by a FLH effect, which puts their validity into question.
@@
As a remedy, the drought threshold can be scaled to account for full-load hour differences between VRE technologies and locations. Threshold scaling relative to different time series attributes can be found in the literature. One option is to scale thresholds relative to their system-specific maximum availability factor [37]. [^6] Technically, in this study, thresholds scale with the maximum generation. Discarding potential curtailment, this is equivalent to scaling with the maximum availability factor.. Focusing on the VRE availability of a single point in time (i.e., the peak point of the duration curves in Figure 4), this approach ignores general availability patterns over the course of a year. Consequently, it does not mitigate differences in full-load hours.
@@
Another option is to use a percentile of the availability time series distribution as threshold [28, 24, 47]. Duration curves are a graphical representation of availability factor distributions (Figure 4). A percentile refers to the availability factor below a certain fraction of the population of time steps falls and depends on the shape of the distribution, which may vary across technologies and regions. Unlike scaling relative to the maximum availability factor, this method recognizes all VRE availability states. However, the percentile approach does not fully adjust for FLH differences.
@@
An effective option for this is to align the drought threshold with each system’s mean availability factor $\overline{avail}$. This method inherently scales with the system’s full-load hours and, therefore, enables comparability across technologies or technology portfolios, regions, and, in the case of comparing multiple years, time [19, 18, 33, 42, 35].
@@
A priori, a meaningful parameterization of the fraction $frac$ of the mean availability factor $\overline{avail}$ for drought identification is not straightforward. Previous studies use various fractions of the mean [27]. While lower thresholds, e.g., $thres=0.1*\overline{avail}$, find very severe but potentially brief and isolated VRE drought events, higher fractions closer to $\overline{avail}$ may result in identifying extended periods of below-average availability, i.e., inter-annual variations in overall VRE availability, rather than true VRE droughts. Relevant fractions likely lie between these extremes. It therefore appears useful to explore a range of thresholds, $thres=frac*\overline{avail}$, for example with $frac$ ranging between 0.1 and 0.9. Obtained events reflect different degrees of low VRE availability, which may be relevant for different types of flexibility options in renewable energy systems.
@@ 2.4 Technological and spatial scope of analysis
The scope of a VRE drought analysis focusing on VRE supply may differ in terms of technological and spatial scope. Each of these dimensions demands specific assumptions, weighting, as well as comparability requirements.
In contrast to single-technology drought analyses, according investigations of a VRE portfolio require a composite time series that combines the availability factors of different VRE technologies. It can be computed as the weighted average of the technology-specific time series. Weighting factors correspond to the technology shares in the assumed overall capacity mix. Considering the temporal and spatial complementarity of wind and solar power [57, 58, 59], such a composite time series allows for within-portfolio balancing, which reflects the generation dynamics of power sectors with high shares of variable renewables. This is in line with several related studies [19, 45, 22, 46]. Another approach excludes or limits such balancing by applying a definition of compound VRE drought events as periods with simultaneously arising drought conditions of multiple VRE technologies [18, 40]. This discards VRE complementarity, hence potentially overestimating droughts characteristics of a VRE portfolio.
For multi-regional settings, drought analyses using VRE availability time series can be based on two extreme scenarios regarding regional interconnection [^7] Here we abstract from within-region grid congestion.: i) perfect interconnection, which abstracts from grid limitations (often referred to as the “copper plate” assumption), and ii) complete isolation, where regions are treated as “island systems” without electricity exchange. When comparing VRE droughts of such island systems, FLH-adjusted threshold scaling is required as discussed above. Realistic patterns of cross-regional electricity exchange fall between these extremes and typically necessitate more sophisticated energy system modeling beyond simple VRE time series analysis.
Similar to the technology portfolio case, multi-regional settings with assumed perfect interconnection require a composite time series that collectively represents all regions. It can be generated as the weighted average of respective regional time series. The weights are assigned based on the relative share of each region’s capacity within the overall system. Unless multiple cross-regional composite time series are to be compared, no FLH-adjustment of the threshold is required. For VRE portfolios across regions, weighting factors need to reflect both within-regional and across-regional shares in the capacity mixes.
@@ 3.1 Positive residual load event definition
In addition to focusing exclusively on VRE supply, research interest in the dynamics between VRE droughts and periods of high electricity demand is intensifying. This becomes more relevant as many countries plan to electrify their heat, transport, and industry sectors, and respective additional loads enter the power sector. Notably during the heating season, electrified heating appliances such as heat pumps increase electricity demand with characteristic profiles [60, 61, 62]. Compound high-demand and VRE drought events, sometimes also referred to as “cold dark doldrums” or, in German, “kalte Dunkelflauten” are emerging as a challenge for decarbonized energy systems that rely on VRE.
@@
PRL events, characterized by high demand and a shortage of VRE supply, indicate the need for firm capacity from thermal generators, or flexibility provided by electricity storage, imports, or demand-side measures [63, 64, 65]. In contrast, negative residual load events represent VRE surplus generation. This surplus can be integrated into the power sector through electricity storage or exports to other areas, or used for electrified loads from coupled sectors, such as heat provision or transport [66]. Alternatively, it can be curtailed [55, 64]. Note that the latter is not an option for PRL periods, which makes them more challenging to deal with than negative residual loads. Accurately quantifying both positive and negative residual load events is relevant for capacity planning of weather-resilient energy systems.
@@
Key PRL characteristics, e.g., duration and energy deficits, can be analyzed using residual load time series $rl_{t}$ (Figure 5). While this can be done analogously to VRE drought analysis, the energy deficit of PRL events can directly be associated with the need for energy that has to be provided by, for instance, long-duration storage. Importantly, the maximum capacity need can be inferred from peak residual load and does not require a dedicated analysis of PRL events.
@@
Unlike VRE drought analysis, there is no need for deliberate threshold parameterization to identify PRL events. Instead, PRL events are simply obtained as periods above zero (gray line in Figure 5). In previous PRL, analysis other approaches such as percentile-based thresholds were used [42]. However, positive residual load directly relates to the need for energy from firm generators, storage, or imports. Hence, from the perspective of a power sector, diverting from thresholds aligned with the zero line appears counter-intuitive.
Figure 5: Exemplary residual load time series and identification methods CAZ, FMAZ, VMAZ, Sequent Peak Algorithm (SPA) (with or without recovery), and SPAadj (upper panel). The green and orange areas indicate the obtained CAZ and FMAZ energy deficit, respectively. The lower panel’s horizontal bars illustrate the varying event duration $dur$ identified by each method.
3.2 Positive residual load event identification methods
@@
We adapt the VRE droughts identification methods discussed in Section 2 to PRL events. Unless stated otherwise, the algorithmic routine is equivalent but recalibrated to identify periods where residual load $rl_{t}$, or its average $MA_{t}^{rl}$, is above zero, rather than below a designated threshold. Analogously to CBT and MBT, we denote Constantly-Above-Zero (CAZ) and Mean-Above-Zero (MAZ) as:
$PRL_{t}^{CAZ}=$$\begin{cases}1&\text{if $rl_{t}>0\ \ \forall\ \ t$}\\ 0&\text{if $rl_{t}\leq 0\ \ \forall\ \ t$}\end{cases}$ $PRL_{t}^{FMAZ}=$$\begin{cases}1&\text{if $MA_{t}^{rl}(intdur)>0\ \ \forall\ \ t$}\\ 0&\text{if $MA_{t}^{rl}(intdur)\leq 0\ \ \forall\ \ t$}\end{cases}$ (5)
@@
Accordingly, the fixed-duration Mean-Above-Zero (FMAZ) and variable-duration Mean-Above-Zero (VMAZ) concepts emulate the FMBT and VMBT approaches.
@@
Analogously to VRE drought analysis, the PRL event duration $dur^{CAZ}$ and $dur^{FMAZ}$ corresponds to the number of consecutive time steps with $PRL_{t}^{CAZ}=1$ and $PRL_{t}^{FMAZ}=1$, respectively. The duration of VMAZ event equals the iteration-specific averaging interval duration $dur^{VMAZ}=intdur^{VMAZ}$.
@@
The VMAZ energy deficit is near zero and hence insignificant. With $t_{k}$ and $t_{l}$ as the start and end points of a PRL event, the CAZ and FMAZ event’s energy deficit (green and orange areas in Figure 5) are defined as:
$ED_{t}^{CAZ}=$ $\sum_{t_{k}}^{t_{l}}rl_{t}$ $ED_{t}^{FMAZ}=$ $\sum_{t_{k}}^{t_{l}}MA_{t}^{rl}$ $\forall\ \ t_{k}\leq t\leq t_{l}$ (6)
@@
The limitations associated with CBT, FMBT, and VMBT discussed in Section 3.2 also extend to their PRL counterparts. The CAZ method, while directly linking event duration, its energy deficit, and its temporal occurrence to for example the required generation and dispatch period of a capacity-constrained firm generator, ignores brief periods of negative residual load adjacent to PRL events (compare the two CAZ events on the left in Figure 5). This can lead to underestimating the need for energy-constrained flexibility options, such as long-duration electricity storage. As for FMAZ, the parameterization of the averaging interval duration $intdur^{MAZ}$ lacks a directly attributable power sector reference and remains arbitrary. As identified FMAZ events may substantially change with varying $intdur^{MAZ}$, they are ambiguous. Further, their interpretation is not straightforward. In general, longer averaging periods may lead to an underestimation of actual PRL challenges as averaging already assumes temporal balancing, for instance, in real-world power systems carried out by electricity storage or other flexibility options. The VMAZ approach, in contrast, obtains unique events. Yet, its relevance remains limited in PRL analysis as it only allows for evaluating PRL event duration but not energy deficit, which converges to zero by design.
@@
The SPA identifies unique PRL while obtaining a meaningful energy deficit. Unlike for VRE droughts, SPA does not require setting an arbitrary threshold for PRL detection, precluding a respective shortcoming in the case of VRE drought identification. SPA allows for brief periods with negative residual load and detects all PRL events within a single iteration. Consequently, we regard it as particularly well-suited for identifying PRL events. Analogously to VRE droughts, its cumulative energy deficit is denoted as:
$ED_{t}^{SPA}=\begin{cases}ED^{SPA}_{t-1}+rl_{t}&\text{if $ED^{SPA}_{t-1}+rl_{t}>0\ \ \forall\ \ t$}\\ 0&\text{if $ED^{SPA}_{t-1}+rl_{t}\leq 0\ \ \forall\ \ t$}\end{cases}$ (7)
@@
The SPA event duration $dur^{SPA}$ refers to the period beginning when the cumulative energy deficit of a single SPA event becomes positive to its global maximum. The SPA method further allows for quantifying a recovery period $recper^{SPA}$. Originating from the hydrology field [67], this period represents the balancing time required for the surplus VRE energy from intermediate and subsequent negative residual load events to offset the energy deficit of a PRL event. It corresponds to the period following the peak cumulative energy deficit $ED_{t}^{SPA}$ to the point in time where this deficit returns to zero $ED_{t}^{SPA}=0$ (Figure 5).
In real-world energy systems, long-duration electricity storage is a key option for dealing with PRL events [22]. Storage always incurs conversion losses, i.e., the amount of electricity recovered from storage is always smaller than previous storage charging. The standard SPA method does not account for such losses. To address this, we introduce an adjusted Sequent Peak Algorithm (SPAadj) method, which incorporates storage efficiency losses arising when balancing positive and negative residual load periods. [^8] For simplicity, we assume that electricity storage does not have any binding capacity constraints with respect to charging, discharging and storage energy. This can be achieved by adjusting all negative residual load values to account for storage round-trip [^9] Technically, this adjustment allots the conversion losses from both charging and discharging to storage charging occurring in intermediate and subsequent negative residual load periods, whose surplus energy is used to offset the energy deficit of a PRL event. efficiency $eff^{LDS}$ as denoted by:
$rl_{t}^{adj}=\frac{rl_{t}}{eff^{LDS}}\ \ \forall\ \ rl_{t}<0$ (8)
In the SPAadj method, the adjusted energy deficit $ED_{t}^{SPA,adj}$ is calculated equivalently to $ED_{t}^{SPA}$. However, assuming that long-duration electricity storage is the relevant flexibility option for dealing with a PRL event, the round-trip efficiency adjustment prolongs the recovery period $recper^{SPA,adj}$ to reflect the time needed to recharge the storage to its pre-PRL state of charge (Figure 5). This approach results in an extended recovery period $recper^{SPA,adj}$ and event duration $dur^{SPA}$ when compared to calculations without considering efficiency losses. This adjustment leads to a more realistic representation of the energy system flexibility need during compound peak demand and VRE drought events. Note that real-world storage options may also incur relevant standing losses, depending on how long the energy is stored. Incorporating these into the SPA concept would be significantly increase complexity and should rather be studied using appropriate energy models.